# Algorithms for Markov Binomial Chains

### v1

@@ Abstract.
We study algorithms to analyze a particular class of Markov population processes that is often used in epidemiology. More specifically, Markov binomial chains are the model that arises from stochastic time-discretizations of classical compartmental models. In this work we formalize this class of Markov population processes and focus on the problem of computing the expected time to termination in a given such model. Our theoretical contributions include proving that Markov binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. We give a PSPACE algorithm for the problem of approximating the time to termination and a direct algorithm for the exact problem in the Blum-Shub-Smale model of computation. Finally, we provide a natural encoding of Markov binomial chains into a common input language for probabilistic model checkers. We implemented the latter encoding and present some initial empirical results showcasing what formal methods can do for practicing epidemilogists.
@@
Work supported by the Flemish inter-university (iBOF) “DESCARTES” project
@@ 1. Introduction
Due to their various important applications, the literature around Markov population models covers interesting mathematical properties of various subclasses. Importantly, and perhaps because of their interest to the systems biology community, there is also a wealth of formal methods to analyze them (see, e.g., [DHSW11, LMW11, BLN18, BBW20, CAdB23]). Importantly, encodings of Markov population models into related stochastic models and language formats accepted by formal-verification tools are known [HJW11].
Most of the literature concerning Markov population processes focuses on continuous-time variants of them. This is in contrast with the class of processes that we study. (Markov) Binomial chains [Bai75] are discrete-time Markov population processes whose transition probabilities are given by a product of probability mass functions of binomial distributions over possible individual transfers between compartments. While this class of model was initially conceived for simple scenarios with small populations, it has recently been used for more complex situations like the analysis of COVID-19 cases in Belgium [AWS+21].
@@
In this work, we initiate the study of algorithms and formal methods for the analysis of binomial chains. We start by formalizing the model in section 2 and then motivating our formalization by giving a self-contained account of how classical compartmental models give rise to binomial chains, in section 3. Then, in section 4, we prove that binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. Finally, in section 5, we give a PSPACE algorithm to approximate the time to termination in a given binomial chain and, in section 6, we also give a direct algorithm for the exact problem (ignoring issues with irrational numbers and the complexity of arithmetic operations). To close the paper, we give an encoding of binomial chains into a common input language for probabilistic model checkers in section 7 and, in section 8, we present some empirical results obtained with an implementation of this encoding.
2. Definitions
@@
2.1. Binomial chains
For the rest of this section, let us fix some $k\in\mathbb{N}$ such that $k\geq 1$.
A binomial chain (BC) is essentially a Markov chain $(S,s_{0},P)$ such that $S\subseteq\mathbb{N}^{k}$ and whose transitions correspond to transfers of individuals between compartments modeled by the components of the state vectors. The probability of each individual transfer is based on a binomial distribution whose success probability is a function of the current state.
  • $\boldsymbol{v}\in\mathbb{N}^{k}$ is the initial state and
  • $\boldsymbol{T}\in\mathbb{L}_{k}^{k\times k}$ is the transfer matrix.
$\forall i\in[k]:\max_{j}M_{ij}$ ${}\leq u_{i}$ (valid binomial outcomes)
and $\boldsymbol{w}$ is the pointwise maximum of $\boldsymbol{0}$ and:
$B(m;n,p)=\binom{n}{m}p^{m}(1-p)^{n-m}$
with the convention that $10^{0}=$ 1.
Remark 1 (Negative populations).
Note that, in general, the second condition from Equation 1 on matrices $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ is necessary to avoid negative populations but it is not sufficient. This is why $\boldsymbol{a}$ is not always $\boldsymbol{0}$, i.e. why the pointwise maximum with $\boldsymbol{0}$ is needed. A natural stronger condition would be to ask that $\boldsymbol{M}\boldsymbol{1}\leq\boldsymbol{u}$. In order to keep the model as simple as possible and to avoid (further) complicating the transition probability expressions, the simpler (insufficient) condition is often preferred (see, e.g., [AWS+21]). In particular, using the stronger condition would require renormalizing Equation 3 since it would exclude certain outcomes of the binomial distributions.
Lemma 2.
Let $\boldsymbol{u},\boldsymbol{w}\in\mathbb{N}^{k}$. We have $P(\boldsymbol{u},\boldsymbol{w})>0$ if and only if there exists $M\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ such that $\max_{j\in[k]}M_{ij}\leq u_{i}$ for all $i\in[k]$, and $M_{ij}>0$ implies $T_{ij}(\boldsymbol{u})>0$ for all $i,j\in[k]$.
2.2. Interesting subclasses
Simple:
Closed:
Acyclic:
Lemma 3.
Proof 2.1.
If all rows of the transfer matrix have at most one nonzero entry, the stronger condition from 1 coincides with the weaker one from our definition of BC. The contrapositive of the converse is also easy to establish: if the BC is not closed then it must be the case that a vector state with a negative component — say, in dimension $i$ — is reached. In turn, from the definition of the binomial distribution, this can only happen if the $i$-th row of $\boldsymbol{T}$ has at least two nonzero entries.
2.3. Computational problems
Termination:
Time to termination:
@@ 3.2. The SIR process
Observe, from (13) and (14), that $S_{t+1}$ and $R_{t+1}$ are independent. Recall that if $X$ and $Y$ are independent then $\Pr(X=x,Y=y)=\Pr(X=x)\Pr(Y=y)=p_{X}(x)p_{Y}(y)$, for $x,y$ in the sample space. Thus, the joint probability mass function $p^{(t+1)}_{S,R}$ of $(S_{t+1},R_{t+1})$ is given by $p^{(t+1)}_{S,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$. Finally, note that $I_{t+1}=N-S_{t+1}-R_{t+1}$, by Lemma 4. Therefore, the joint probability mass function $p^{(t+1)}_{S,I,R}$ of $(S_{t+1},I_{t+1},R_{t+1})$ is also given by $p^{(t+1)}_{S,I,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$.
@@ 3.3. A discrete-time Markov chain induced by the binomial chain
Finally, the transfer matrix $\boldsymbol{T}$ of the BC whose induced Markov chain we defined above, can be inferred from Figure 1 and Equation 15. For concreteness, we give it below.
$\boldsymbol{T}=\begin{pmatrix}0&\boldsymbol{m}\mapsto h\beta m_{2}&0\\ 0&0&\boldsymbol{m}\mapsto h\gamma\\ 0&0&0\end{pmatrix}$
Note that the binomial chain is both closed and acyclic.
@@ 4.4. Acyclic binomial chains are absorbing
We will now define a second sequence of states $\boldsymbol{w^{(1)}}\dots\boldsymbol{w^{(k)}}$, such that $\boldsymbol{w^{(1)}}=\boldsymbol{u}$ and $\boldsymbol{w^{(k)}}=\boldsymbol{w}$, with their corresponding matrices $\boldsymbol{N^{(1)}},\dots,\boldsymbol{N^{(k-1)}}$. It will be clear that $\boldsymbol{N^{(\ell)}}\in\mathbf{Wit}(\boldsymbol{w^{(\ell)}},\boldsymbol{w^{(\ell+1)}})$, for all $1\leq\ell<k$, based on the triangularity of the $\boldsymbol{M^{(\ell)}}$ and how we will define our $\boldsymbol{N^{(\ell)}}$. Afterwards, we will argue that the corresponding products of the binomial probability mass functions are positive. The intuition is that $\boldsymbol{N^{(\ell)}}$ will correspond to the sum of the $\ell$-th rows of all $\boldsymbol{M^{(\ell^{\prime})}}$ with $\ell^{\prime}>\ell$ and all the rows of $\boldsymbol{M^{(\ell)}}$ strictly below the $\ell$-th one.
@@
$\boldsymbol{N^{(\ell)}_{i}}=\begin{cases}\boldsymbol{0}&\text{if }i<\ell\\ \sum_{\ell^{\prime}\geq\ell}\boldsymbol{M^{(\ell^{\prime})}_{i}}&\text{if }i=\ell\\ \boldsymbol{M^{(\ell)}_{i}}&\text{otherwise.}\end{cases}$
@@
Let $\ell\in[k]$ be arbitrary. By triangularity of the original witness matrices $\boldsymbol{M^{(\ell^{\prime})}}$ and the definition of the new ones $\boldsymbol{N^{(\ell^{\prime})}}$, we get that $w^{(\ell)}_{i}=w_{i}$ for all $i\leq\ell$ and $w^{(\ell)}_{i}\geq u^{(\ell)}_{i}$ for all $i>\ell$. In particular, this means that $\mathrm{supp}(\boldsymbol{w^{(\ell)}})=\mathrm{supp}(\boldsymbol{u^{(\ell)}})$. Recall that every $T_{ij}$ can be written as $\boldsymbol{x}\mapsto\boldsymbol{a}^{\intercal}\boldsymbol{x}+b$ with $\boldsymbol{a}$ and $b$ being nonnegative. Hence, if $T_{ij}(\boldsymbol{x})>0$ and $\mathrm{supp}(\boldsymbol{y})=\mathrm{supp}(\boldsymbol{x})$, for vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{N}^{k}$, then $T_{ij}(\boldsymbol{y})>0$. We thus conclude that for all $1\leq\ell^{\prime}<\ell$ the following holds.
@@ 6. Time to termination for SIR models in the Blum-Shub-Smale model
Let us write $Q$ for the set of states of the Markov chain induced by the given SIR binomial chain $(\boldsymbol{v},\boldsymbol{T})$, so $Q\subseteq\{0,1,\dots,N\}^{3}$, where $N=\left\lVert\boldsymbol{v}\right\rVert_{1}$, and the components of a state $\boldsymbol{m}\in Q$ correspond to susceptible, infectious, and recovered, in that order. We also write $A\subseteq Q$ for its absorbing states, thus $A=\{\boldsymbol{m}\in Q\mid m_{2}=0\}$ by 5. Note that, since SIR binomial chains are closed, one of the components of the vector states is redundant and $|Q|\leq N^{2}$. Hence, the dimensions of $\boldsymbol{P}$ (the matrix representation of the transition probability function of the induced Markov chain) are at most $N^{2}$.

### v2

@@ Abstract.
We study algorithms to analyze a particular class of Markov population processes that is often used in epidemiology. More specifically, Markov binomial chains are the model that arises from stochastic time-discretizations of classical compartmental models. In this work we formalize this class of Markov population processes and focus on the problem of computing the expected time to termination in a given such model. Our theoretical contributions include proving that Markov binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. We give a PSPACE algorithm for the problem of approximating the time to termination and a direct algorithm for the exact problem in the Blum-Shub-Smale model of computation. Finally, we provide a natural encoding of Markov binomial chains into a common input language for probabilistic model checkers. We implemented the latter encoding and present some initial empirical results showcasing what formal methods can do for practicing epidemiologists.
@@
This work was supported by the Flemish inter-university (iBOF) “DESCARTES” project.
@@ 1. Introduction
Due to their various important applications, the literature around Markov population models covers interesting mathematical properties of various subclasses. Importantly, and perhaps because of their interest to the systems biology community, there is also a wealth of formal methods to analyze them (see, e.g., [DHSW11, LMW11, BLN18, BBW20, CAdB23]). Importantly, encodings of Markov population models into related stochastic models and language formats accepted by formal-verification tools are known [HJW11]. In addition, techniques from fluid limits can be used to efficiently infer stochastic information of interest from some of the continuous-time classes of such models [BH12, Bor10].
Most of the literature concerning Markov population processes focuses on continuous-time variants of them. This is in contrast with the class of processes that we study. (Markov) Binomial chains [Bai75] are discrete-time Markov population processes whose transition probabilities are given by a product of probability mass functions of binomial distributions over possible individual transfers between compartments. While this class of model was initially conceived for simple scenarios with small populations, it has recently been used for more complex situations like the analysis of COVID-19 cases [AWS+21, PWC+22]. Recently, it has also been established that (for epidemiological tasks) discrete-time models are as general and as flexible as their continuous-time counterparts, yet they are simpler to parameterize on the basis of data and to implement computationally [DOPB21].
@@
In this work, we initiate the study of algorithms and formal methods for the analysis of binomial chains. We start by giving a self-contained account of how classical compartmental models give rise to binomial chains, in section 3. The derivation of a binomial chain from compartmental models in that section is meant as a way to motivate the class of models and to provide useful epidemiological context and intuition to the unfamiliar reader. Importantly, we do not mean for the translation to be used as a way to analyze compartmental models by approximating them with a binomial chain. As per the motivation above, our target use case is when the ground-truth model is already a binomial chain fitted to historical data by epidemiologists. In in section 4, we then move to formalizing the general model of binomial chains. After that, in section 5, we prove that binomial chains whose flow of individuals through compartments is acyclic almost surely terminate. Finally, in section 6, we give a PSPACE algorithm to approximate the time to termination in a given binomial chain and, in section 7, we also give a direct algorithm for the exact problem (ignoring issues with irrational numbers and the complexity of arithmetic operations). To close the paper, we give an encoding of binomial chains into a common input language for probabilistic model checkers in section 8 and, in section 9, we present some empirical results obtained with an implementation of this encoding.
2. Preliminaries
@@ 3.2. The SIR process
Observe, from (10) and (11), that $S_{t+1}$ and $R_{t+1}$ are independent. That is, the events $S_{t+1}=n_{1}$ and $R_{t+1}=n_{3}$ are independent if we condition on $(S_{t},I_{t},R_{t})=(m_{1},m_{2},m_{3})$. Since $I_{t+1}=N-S_{t+1}-R_{t+1}$, by 1, the SIR process satisfies the Markov property. It remains to compute the exact probability of such a transition.
Recall that if $X$ and $Y$ are independent then:
$\Pr(X=x,Y=y)=\Pr(X=x)\Pr(Y=y)=p_{X}(x)p_{Y}(y),$
for $x,y$ in the sample space. Thus, the joint probability mass function $p^{(t+1)}_{S,R}$ of $(S_{t+1},R_{t+1})$ is given by $p^{(t+1)}_{S,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$. According to 1, $I_{t+1}=N-S_{t+1}-R_{t+1}$, so the joint probability mass function $p^{(t+1)}_{S,I,R}$ of $(S_{t+1},I_{t+1},R_{t+1})$ is also given by $p^{(t+1)}_{S,I,R}=p^{(t+1)}_{S}p^{(t+1)}_{R}$.
@@ 3.3. A discrete-time Markov chain induced by the binomial chain
4. Binomial chains
Let $k\in\mathbb{N}$ such that $k\geq 1$. A binomial chain (BC) is essentially a Markov chain $(S,s_{0},P)$ such that $S\subseteq\mathbb{N}^{k}$ and whose transitions correspond to transfers of individuals between compartments modeled by the components of the state vectors. The probability of each individual transfer is based on a binomial distribution whose success probability is a function of the current state.
  • $\boldsymbol{v}\in\mathbb{N}^{k}$ is the initial state and
  • $\boldsymbol{T}\in\mathbb{L}_{k}^{k\times k}$ is the transfer matrix.
{exa}
The transfer matrix $\boldsymbol{T}$ of the SIR binomial chain from section 3 is given below.
$\boldsymbol{T}=\begin{pmatrix}0&\boldsymbol{m}\mapsto h\beta m_{2}&0\\ 0&0&\boldsymbol{m}\mapsto h\gamma\\ 0&0&0\end{pmatrix}$ (14)
It can be inferred from Figure 1 and Equation 12. (A formal description of how this is done follows, here we are just interested in conveying the intuition of the model.) Notice that the entries in the matrix correspond to the arguments of the exponentials in the expression that gives the transition probability.$\triangle$
{exa}
The adjacency matrix $\left.\boldsymbol{T}\right|_{\mathbb{B}}$ of the SIR binomial chain from section 3 is given below.
$\left.\boldsymbol{T}\right|_{\mathbb{B}}=\begin{pmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{pmatrix}$ (15)
This one can be obtained directly from Figure 1. Indeed, it is nothing more than the adjacency matrix of that graph showing what transfers ae possible between compartments. $\triangle$
$\min_{i\in[k]}\max_{j\in[k]}M_{ij}$ ${}\leq u_{i}$ (valid binomial outcomes)
and $\boldsymbol{w}$ is the pointwise maximum of $\boldsymbol{0}$ and:
{exa}
Let us turn once more to the SIR binomial chain from section 3. Recall that $k=3$. Further consider states $\boldsymbol{u}=(10,3,2)^{\intercal}$ and $\boldsymbol{w}=(8,2,5)^{\intercal}$. We use matrices $\boldsymbol{M}\in\mathbb{N}^{k\times k}$ to encode information about transfers between compartments. To move from state $\boldsymbol{u}$ to $\boldsymbol{w}$ by transferring individuals along edges from Figure 1, thus nonzero entries from Equation (15), we can use the following matrix.
$\boldsymbol{M}=\begin{pmatrix}0&2&0\\ 0&0&3\\ 0&0&0\end{pmatrix}$
Moving $2$ susceptible individuals to the infectious compartment, and $3$ infectious ones to the recovered compartment, yields $\boldsymbol{w}$ if we start from $\boldsymbol{u}$. In symbols, this is just $\boldsymbol{w}=\boldsymbol{u}+(\boldsymbol{1}^{\intercal}\boldsymbol{M})^{\intercal}-\boldsymbol{M}\boldsymbol{1}$. In this example, it turns out that $\boldsymbol{M}$ is the unique matrix that satisfies that equation and the constraints imposed above. Furthermore, the pointwise maximum with $\boldsymbol{0}$ is unnecessary. However, uniqueness and positivity of $\boldsymbol{u}+(\boldsymbol{1}^{\intercal}\boldsymbol{M})^{\intercal}-\boldsymbol{M}\boldsymbol{1}$ are not guaranteed in general.$\triangle$
$B(m;n,p)=\binom{n}{m}p^{m}(1-p)^{n-m}$
with the convention that $10^{0}=$ 1.
{exa}
We continue with the situation from section 4. Recall that we argued $\boldsymbol{M}$ was the unique matrix satisfying the imposed constraints. That is, $\{\boldsymbol{M}\}=\mathbf{Wit}(\boldsymbol{u},\boldsymbol{v})$. Now, from Equation 12 we know that:
$P(\boldsymbol{u},\boldsymbol{w})={}$ $\binom{10}{2}(\exp(-h\beta 3))^{8}(1-\exp(-h\beta 3)^{2}$
${}\binom{3}{3}(1-\exp(-h\gamma))^{3}\exp(-h\gamma)^{0}$
and this coincides with what we obtain from Equation (18) using Equation (14) for the transfer matrix. Namely, we get:
$P(\boldsymbol{u},\boldsymbol{w})=B(2;10,1-\exp(h\beta 3))B(3;3,1-\exp(h\gamma))$
as expected.$\triangle$
Remark 3 (Negative populations).
Note that, in general, the second condition from Equation 16 on matrices $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ is necessary to avoid negative populations, but it is not sufficient. This is why the pointwise maximum with $\boldsymbol{0}$, from Equation 17, is needed. A natural stronger condition would be to ask that $\boldsymbol{M}\boldsymbol{1}\leq\boldsymbol{u}$. In order to keep the model as simple as possible and to avoid (further) complicating the transition probability expressions, the simpler (insufficient) condition is often preferred (see, e.g., [AWS+21]). In particular, using the stronger condition would require renormalizing Equation 18 since it would exclude certain outcomes of the binomial distributions.
Lemma 4.
Let $\boldsymbol{u},\boldsymbol{w}\in\mathbb{N}^{k}$. We have $P(\boldsymbol{u},\boldsymbol{w})>0$ if and only if there exists $\boldsymbol{M}\in\mathbf{Wit}(\boldsymbol{u},\boldsymbol{w})$ such that $M_{ij}>0$ implies $T_{ij}(\boldsymbol{u})>0$ for all $i,j\in[k]$.
4.1. Interesting subclasses
Simple:
Closed:
Acyclic:
Lemma 5.
Proof 4.1.
If all rows of the transfer matrix have at most one nonzero entry, the stronger condition from 3 coincides with the weaker one from our definition of BC. The contrapositive of the converse is also easy to establish: if the BC is not closed then it must be the case that a vector state with a negative component — say, in dimension $i$ — is reached. In turn, from the definition of the binomial distribution, this can only happen if the $i$-th row of $\boldsymbol{T}$ has at least two nonzero entries.
4.2. Computational problems
Termination:
Time to termination:
@@ 5.4. Acyclic binomial chains are absorbing
Finally, we write $\boldsymbol{M_{i}^{(\ell)}}$ to denote the $i$-th row $(M^{(\ell)}_{i1},M^{(\ell)}_{i2},\dots,M^{(\ell)}_{ik})$ of $\boldsymbol{M^{(\ell)}}$.
We will now define a second sequence of states $\boldsymbol{w^{(1)}}\dots\boldsymbol{w^{(k)}}$, such that $\boldsymbol{w^{(1)}}=\boldsymbol{u^{(1)}}$ and $\boldsymbol{w^{(k)}}=\boldsymbol{u^{(n)}}$, with their corresponding matrices $\boldsymbol{N^{(1)}},\dots,\boldsymbol{N^{(k-1)}}$. It will be clear that $\boldsymbol{N^{(\ell)}}\in\mathbf{Wit}(\boldsymbol{w^{(\ell)}},\boldsymbol{w^{(\ell+1)}})$, for all $1\leq\ell<k$, based on the triangularity of the $\boldsymbol{M^{(\ell)}}$ and how we will define our $\boldsymbol{N^{(\ell)}}$. Afterwards, we will argue that the corresponding products of the binomial probability mass functions are positive. The intuition is that $\boldsymbol{N_{i}^{(\ell)}}$ groups all transfers encoded by the $\boldsymbol{M^{(\ell^{\prime})}}$, for $1\leq\ell^{\prime}<n$, from the $i$-th compartment to other compartments $j\geq i$.
@@
$\boldsymbol{N^{(\ell)}_{i}}=\begin{cases}\sum_{1\leq\ell^{\prime}\leq n}\boldsymbol{M^{(\ell^{\prime})}_{i}}&\text{if }i=\ell\\ \boldsymbol{0}&\text{otherwise.}\end{cases}$
@@
Let $\ell\in[k]$ be arbitrary. By triangularity of the original witness matrices $\boldsymbol{M^{(\ell^{\prime})}}$ and the definition of the new ones $\boldsymbol{N^{(\ell^{\prime})}}$, we get that $w^{(\ell)}_{i}=u^{(n)}_{i}$ for all $i<\ell$ and $w^{(\ell)}_{i}\geq u^{(1)}_{i}$ for all $i\geq\ell$. In particular, since all $\boldsymbol{u^{(\ell^{\prime})}}$ have the same support, this means that:
$\mathrm{supp}(\boldsymbol{w^{(\ell)}})=\mathrm{supp}(\boldsymbol{u^{(n)}})=\mathrm{supp}(\boldsymbol{u^{(1)}})=\mathrm{supp}(\boldsymbol{u^{(\ell)}}).$
Recall that every $T_{ij}$ can be written as $\boldsymbol{x}\mapsto\boldsymbol{a}^{\intercal}\boldsymbol{x}+b$ with $\boldsymbol{a}$ and $b$ being nonnegative. Hence, if $T_{ij}(\boldsymbol{x})>0$ and $\mathrm{supp}(\boldsymbol{y})=\mathrm{supp}(\boldsymbol{x})$, for vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{N}^{k}$, then $T_{ij}(\boldsymbol{y})>0$. We thus conclude that for all $1\leq\ell^{\prime}<\ell$ the following holds.
@@ 7. Time to termination for SIR models in the Blum-Shub-Smale model
Let us write $Q$ for the set of states of the Markov chain induced by the given SIR binomial chain $(\boldsymbol{v},\boldsymbol{T})$, so $Q\subseteq\{0,1,\dots,N\}^{3}$, where $N=\left\lVert\boldsymbol{v}\right\rVert_{1}$, and the components of a state $\boldsymbol{m}\in Q$ correspond to susceptible, infectious, and recovered, in that order. We also write $A\subseteq Q$ for its absorbing states, thus $A=\{\boldsymbol{m}\in Q\mid m_{2}=0\}$ by 2. Note that, since SIR binomial chains are closed, one of the components of the vector states is redundant and $|Q|\leq N^{2}$. Hence, the dimensions of $\boldsymbol{P}$ (the matrix representation of the transition probability function of the induced Markov chain) are at most $N^{2}$.